# Implicit Safe Set Algorithm for Provably Safe Reinforcement Learning

### v1

@@ 7.3.2 Proof the Theorem 2
Figure 5: The illustration of forward invariance under the constraint of $\phi(x_{t+1})\leq\max\{\phi(x_{t})-\eta,0\})-\sigma^{*}$.
@@
2.2) $d_{{t_{0}}_{i}}>d_{min}+\sigma^{*}$. Note that $x_{t}\in\mathcal{S}$ and $d_{t_{i}}>d_{min}+\sigma^{*}$ indicate $\sigma+d^{n}_{min}-d_{t_{i}}^{n}-k\dot{d_{t_{i}}}\leq 0$ and $d_{min}-d_{t_{i}}<0$, whereas $\dot{d}$ can either $<0$ or $\geq 0$.
@@ 8 Experimental Results
    How does the discrepancy between continuous-time system safe control and discrete-time system safe control affect the evolution of the safety index in discrete-time system? (Answered in Section 8.1)
@@
8.1 Toy Problem Experiment Details
Figure 6: The toy experimental platform. The blue circle is a static obstacle, and red circle is the robot, whose heading direction is illustrated as the black arrow.
To demonstrate the discrepancy between continuous-time systems and discrete-time systems, we build a toy problem environment to show that directly solving (2) in discrete-time systems will lead to safety violations. The toy robot is a 3-state unicycle model with state $x=[p_{x},p_{y},\theta]$, where $p_{x},p_{y}$ denote the coordinates on x-axis and y-axis, respectively, and $\theta$ is the heading direction. The control inputs of the toy environment as $[v,w]$, where $v$ denotes the velocity of the robot, and $w$ denotes the angular velocity of the robot. Next, we define the underlying dynamics of the robot for toy environment in control affine form as following:
$\begin{split}x_{t+1}&=x_{t}+\dot{x}_{t}dt=x_{t}+\begin{bmatrix}\cos{(\theta)}dt&0\\ \sin{(\theta)}dt&0\\ 0&dt\end{bmatrix}u\end{split}$ (60)
8.1.1 Experimental Setup: Toy Problem
@@
Therefore, in Figure 6 we only consider one obstacle. For simplicity consideration, we currently assume the obstacle is static, which can be easily extended to moving obstacles.
Next, we define the safety index $\phi(x)$ by constraining the perpendicular line from the center of the obstacle to tne robot heading direction to be larger than $(R+r)$, which is shown as follows:
$\phi(x)=(r+R)^{2}-(({p_{x}}_{0}-p_{x})\sin(\theta)-({p_{y}}_{0}-p_{y})\cos(\theta))^{2}$ (61)
where $r,R$ are the radius of robot and obstacle, respectively. $[{p_{x}}_{0},{p_{y}}_{0}]$ is the location of the obstacle. It can be easily shown that $\phi(x)<0$ will ensure the robot is collision-free with environmental obstacles. Finally, we design a dummy nominal policy to control the Point robot to move forward with constant velocity, which has no safety guarantee. Therefore, if the initial system state is unsafe, our proposed methods are expected to generate the safe control to gradually drive the system state to the safe set where $\phi(x)<0$, and remains in the safe set.
8.1.2 Results: Toy Problem
Figure 7: Safety Index evolution curves comparison.
@@ 8.1.2 Results: Toy Problem
8.2 Safety Gym Experiment Details
To demonstrate the effectiveness of the proposed implicit safe set algorithms, we conduct evaluation experiments on the safe reinforcement learning benchmark environment Safety Gym (?, ?).
@@ 8.2 Safety Gym Experiment Details
Figure 8: The environmental settings for benchmark problems in Safety Gym.
@@
Figure 9: The PPO-ISSA structure.
The methods in the comparison group include: unconstrained RL algorithm PPO (?) and constrained safe RL algorithms PPO-Lagrangian, CPO (?) and PPO-SL (PPO-Safety Layer) (?). We select PPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. We set the limit of cost to 0 for both PPO-Lagrangian and CPO since we aim to avoid any violation of the constraints. To make sure ISSA can complete tasks while guaranteeing safety, we use a PPO agent as the nominal policy and ISSA as a safety layer to solve (5), we call this structure as PPO-ISSA, and it is illustrated in Figure 9. Such safety layer structure has also been used in PPO-SL (?) which leverages offline dataset to learn a linear safety-signal model and then construct a safety layer via analytical optimization. For all experiments, we use neural network policies with separate feedforward MLP policy and value networks of size (256, 256) with tanh activations. More details are as follows.
8.2.1 Environment Settings
@@
The underlying dynamics of Safety Gym is directly handled by MuJoCo physics simulator (?). This indicates the dynamics is not explicitly accessible but rather can be implicitly evaluated, which is suitable for our proposed implicit safe set algorithm. The implementation of block-box dynamics for ISSA is through simulation in the MuJoCo physics simulator and recovering to the pre-simulated state.
@@
The state space is composed of various physical quantities from standard robot sensors (accelerometer, gyroscope, magnetometer, and velocimeter) and lidar (where each lidar sensor perceives objects of a single kind). The state spaces of all the test suites are summarized in Table 1. Note that Vase is another type of constraint in Safety Gym (?) and all the returns of vase lidar are zero vectors (i.e., $[0,0,\cdots,0]\in\mathbb{R}^{16}$) in our experiments since none of our eight test suites environments have vases.
Table 1: The state space components of different test suites environments.
@@
For all the experiments, the control space $\mathcal{U}\subset\mathbb{R}^{2}$. The first dimension $u_{1}\in[-10,10]$ is the control space of moving actuator, and second dimension $u_{2}\in[-10,10]$ is the control space of turning actuator.
8.2.2 Policy Settings
[…]
Table 2: Important hyper-parameters of PPO, PPO-Lagrangian, CPO, PPO-SL and PPO-ISSA
@@ 8.2.2 Policy Settings
8.2.3 Safety Index Settings
@@
Table 3: Experiment-specific parameters of safety index design for PPO-ISSA.
@@
8.3 Evaluating PPO-ISSA and Comparison Analysis
@@
Figure 10: Cost changes over 100 time steps of PPO and PPO-ISSA-CTrigger starting from the same unsafe state over 20 trails.
PPO-Lagrangian and CPO fail since both methods rely on trial-and-error to enforce constraints while ISSA is able to guarantee forward invariance by Theorem 1. We also observe that PPO-SL fails to lower the violation during training, due to the fact that the linear approximation of cost function $c(x_{t+1})\approx c(x_{t})+g(x_{t},w)^{T}u$ (?) becomes inaccurate when the dynamics are highly nonlinear like the ones we used in MuJoCo (?). More importantly, PPO-SL cannot guarantee that these always exist a feasible safe control to lower the cost, since they directly use the user defined cost function which cannot always ensure feasibility. More detailed metrics for comparison and experimental results on test suites with 1 constraint are summarized in Section 8.3.1.
8.3.1 Metrics Comparison
In this section, we report all the results of eight test suites by three metrics defined in Safety Gym (?):
@@ 8.3.1 Metrics Comparison
The average episode return $J_{r}$ and the average episodic sum of costs $M_{c}$ were obtained by averaging over the last five epochs of training to reduce noise. Cost rate $\rho_{c}$ was just taken from the final epoch. We report the results of these three metrics in Table 4 normalized by PPO results. We calculate the converged reward $\bar{J}_{r}$ percentage of PPO-ISSA compared to other three safe RL baseline methods (PPO-Lagrangian, CPO and PPO-SL) over eight control suites. The computed mean reward percentage is $95\%$, and the standard deviation is $9\%$. Therefore we conclude that PPO-ISSA is able to gain $95\%\pm 9\%$ cumulative reward compared to state-of-the-art safe DRL methods.
@@
Table 4: Normalized metrics obtained from the policies at the end of the training process, which is averaged over eight test suits environments and five random seeds.
@@
Figure 12: Average episodic return, episodic cost and overall cost rate of constraints of PPO-ISSA and baseline methods on 4-constraint environments over five seeds.
@@
Figure 14: Average performance of PPO-ISSA and baseline methods on 1-constraint environments over five random seeds. The three rows represent average episodic return, average episodic cost and overall cost rate of constraints. The safety advantage of safe RL baseline methods over unconstrained RL method (PPO) becomes trivial as the constraint number and constraint size decrease, where the cost rate and average episode cost of PPO-Lagrangian, CPO and PPO are nearly the same when there is only one constraint with size 0.05.
To validate the finite time convergence of Theorem 2 that ISSA combined with CTrigger algorithm (PPO-ISSA-CTrigger) can ensure the finite time convergence to $\mathcal{X}_{S}$ for discrete-time system, we further compare the cost evolution of PPO and PPO-ISSA-CTrigger agents when starting from the same unsafe state (i.e., cost $>0$). The comparison results are shown in Figure 10, which shows that PPO-ISSA-CTrigger can converge to $\mathcal{X}_{S}$ within 100 time steps across all experiments while cost evolution of PPO agents fluctuates wildly without preference to converge to safe set. The cost changes of PPO-ISSA-CTrigger aligns with ours theory of finite time convergence.
Figure 15: Average return of PPO-ISSA with different safety index design on Goal-Hazard4-0.15.
8.4 Feasibility of Safety Index Synthesis
@@
Figure 16: Heat maps of the difference of safety index $\Delta\phi=\phi(f(x,u))-\phi(x)$. The x-axis $u_{1}$ represents the control space of moving actuator, and the y-axis $u_{2}$ represents the control space of turning actuator.
@@
8.5 Ablation Study
8.5.1 Sensitivity Analysis
@@
Table 5: Normalized computation time and return under different number of vectors in ISSA. These results are average on 100 ISSA runs over five random seeds on Goal-Hazard4-0.15.
@@
Figure 17: Average episodic return and episodic cost of PPO-ISSA and baseline methods on Goal-hazard1-0.15 environment of a doggo robot over five seeds.
@@ 8.5.2 Scalability Analysis
Figure 18: doggo robot: a quadrupedal robot with bilateral symmetry with 12-dimensional control space.
@@
Figure 19: Average performance of PPO-ISSA and baseline methods on Goal-hazard1-0.15 environment of doggo robot over five seeds. The three columns represent average episode return, average episode cost and overallcost rate of constraints.
8.5.3 ISSA in higher dimensional constrol systems
We test ISSA with a doggo robot as shown in Figure 18 with 12 dimensional control space and 80 dimensional state space. We evaluate ISSA with the doggo robot in the Goal-Hazard1-0.15 suite. As shown in Figure 19, ISSA is able to guarantee zero-violation in higher dimensional control systems. We notice that in system with higher control dimensions, safe RL methods like PPO-Lagrangian and CPO perform poorly compared to PPO, which demonstrates the constrained RL algorithms struggle to learn good reward performance for complex locomotion behavior. Similar comparison results are also reported in Safety Gym benchmarks (?). In contrast, PPO-ISSA is able to achieve the best reward performance while guaranteeing zero safety violation, showing the scalability of ISSA to achieve satisfying reward performance in systems with higher control dimensions.
@@ 8.5.3 ISSA in higher dimensional constrol systems
Table 6: Average Computation time, succuss ratio of ISSA phase 1 and number of safe control candidates of 200 ISSA runs on Goal-Hazard1-0.15.
@@
8.5.4 AdamBA in higher dimensional space
@@
Figure 20: The average computation time of non-parallel AdamBA and parallel AdamBA on toy problem with different control dimensions and number of vectors.
9 Conclusion and Future Work
@@
There are two major directions for future work. Firstly, we will further generalize the safety index synthesis rule to cover a wider range of applications other than collision avoidance in 2D. Secondly, we will further speed up the implicit model evaluation step by replacing the physical engine based simulator with a learned surrogate model while taking the learned dynamics error into account.
Acknowledgments
This work is supported by Amazon Research Award.

### v2

@@ 7.3.2. Proof the Theorem 1
Figure 6. The illustration of forward invariance under the constraint of $\phi(x_{t+1})\leq\max\{\phi(x_{t})-\eta,0\})-\sigma^{*}$.
@@
2.2) $d_{{t_{0}}_{i}}>d_{min}+\sigma^{*}$. Note that $x_{t}\in\mathcal{S}$ and $d_{t_{i}}>d_{min}+\sigma^{*}$ indicate $\sigma+d^{n}_{min}-d_{t_{i}}^{n}-k\dot{d_{t_{i}}}\leq 0$ and $d_{min}-d_{t_{i}}<0$, whereas $\dot{d}$ can either <0<0 or $\geq 0$.
@@ 8. Experimental Results
    How does the discrepancy between continuous-time safe control and discrete-time safe control affect the evolution of the safety index in discrete-time systems? (Answered in Section 8.1)
@@
8.1. Toy Problem Experiment Details
To demonstrate the discrepancy between continuous-time systems and discrete-time systems, we build a toy problem environment to show that directly applying the result from (2) in discrete-time systems will lead to safety violations. The toy robot is a 3-state unicycle model with state $x=[p_{x},p_{y},\theta]$, where $p_{x},p_{y}$ denote the coordinates on the x-axis and y-axis, respectively, and $\theta$ is the heading direction. The control inputs of the toy environment are $[v,w]$, where $v$ denotes the velocity of the robot, and $w$ denotes the angular velocity of the robot. Next, we define the underlying dynamics of the robot for the toy environment in control affine form as follows:
(60) $\begin{split}x_{t+1}&=x_{t}+\dot{x}_{t}dt=x_{t}+\begin{bmatrix}\cos{(\theta)}dt&0\\ \sin{(\theta)}dt&0\\ 0&dt\end{bmatrix}u\end{split}$
Figure 7. The toy experimental platform. The blue circle is a static obstacle, and red circle is the robot, whose heading direction is illustrated as the black arrow.
8.1.1. Experimental Setup: Toy Problem
@@
Therefore, in Figure 7 we only consider one obstacle. For simplicity, we currently assume the obstacle is static, which can be easily extended to moving obstacles.
Next, we define the safety index $\phi(x)$ by constraining the perpendicular line from the center of the obstacle to the robot heading direction to be larger than $(R+r)$, which is shown as follows:
(61) $\phi(x)=(r+R)^{2}-(({p_{x}}_{0}-p_{x})\sin(\theta)-({p_{y}}_{0}-p_{y})\cos(\theta))^{2}$
where $r,R$ are the radii of the robot and obstacle, respectively. ${p_{x}}_{0},{p_{y}}_{0}]$ is the location of the obstacle. It can be easily shown that $\phi(x)<0$ will ensure the robot is collision-free with environmental obstacles. Finally, we design a dummy nominal policy to control the point robot to move forward with constant velocity, which has no safety guarantee. Therefore, if the initial system state is unsafe, our proposed methods are expected to generate the safe control to gradually drive the system state to the safe set where $\phi(x)<0$, and remain in the safe set.
8.1.2. Results: Toy Problem
@@
Figure 8. Safety Index evolution curves comparison.
@@
8.2. Safety Gym Experiment Details
In recent years, numerous benchmark environments for safe control have emerged (ray2019benchmarking, 32, 20, 46, 35). We adopt Safety Gym (ray2019benchmarking, 32) as our testing platform to evaluate the effectiveness of the proposed implicit safe set algorithms.
@@ 8.2. Safety Gym Experiment Details
Figure 9. The environmental settings for benchmark problems in Safety Gym.
@@
The methods in the comparison group include: unconstrained RL algorithm PPO (schulman2017proximal, 34) and constrained safe RL algorithms PPO-Lagrangian, CPO (achiam2017constrained, 2) and PPO-SL (PPO-Safety Layer) (dalal2018safe, 10). We select PPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. We set the limit of cost to 0 for both PPO-Lagrangian and CPO since we aim to avoid any violation of the constraints. To make sure ISSA can complete tasks while guaranteeing safety, we use a PPO agent as the nominal policy and ISSA as a safety layer to solve (5), we call this structure as PPO-ISSA, and it is illustrated in Figure 10. Such safety layer structure has also been used in PPO-SL (dalal2018safe, 10) which leverages offline dataset to learn a linear safety-signal model and then construct a safety layer via analytical optimization.
Figure 10. The PPO-ISSA structure.
For all experiments, we use neural network policies with separate feedforward MLP policy and value networks of size (256, 256) with tanh activations. More details are as follows.
8.2.1. Environment Settings
@@
The underlying dynamics of Safety Gym is directly handled by MuJoCo physics simulator (todorov2012mujoco, 37). This indicates the dynamics is not explicitly accessible but rather can be implicitly evaluated, which is suitable for our proposed implicit safe set algorithm. The implementation of block-box dynamics for ISSA is through simulation in the MuJoCo physics simulator and recovering to the pre-simulated state.
@@
The state space is composed of various physical quantities from standard robot sensors (accelerometer, gyroscope, magnetometer, and velocimeter) and lidar (where each lidar sensor perceives objects of a single kind). The state spaces of all the test suites are summarized in Table 1. Note that Vase is another type of constraint in Safety Gym (ray2019benchmarking, 32) and all the returns of vase lidar are zero vectors (i.e., $0,0,\cdots,0]\in\mathbb{R}^{16}$) in our experiments since none of our eight test suites environments have vases.
Table 1. The state space components of different test suites environments.
@@
For all the experiments, the control space $\mathcal{U}\subset\mathbb{R}^{2}$. The first dimension $u_{1}\in[-10,10]$ is the control space of moving actuator, and second dimension $u_{2}\in[-10,10]$ is the control space of turning actuator. For each actuator the maximum torque corresponds to $a_{\max}=2.5\,\text{m/s}^{2}$ and $\omega_{\max}=4.0\,\text{rad/s}$; within these limits the inner-loop PID reaches the desired $(a,w)$ in $\leq 6\,\text{ms}$, satisfying Assumption 1.
8.2.2. Policy Settings
Detailed parameter settings are shown in Table 2. All the policies in our experiments use the default hyper-parameter settings hand-tuned by Safety Gym (ray2019benchmarking, 32) except the $\text{cost limit}=0$ for PPO-Lagrangian and CPO.
Table 2. Important hyper-parameters of PPO, PPO-Lagrangian, CPO, PPO-SL and PPO-ISSA
@@ 8.2.2. Policy Settings
8.2.3. Safety Index Settings
@@
Table 3. Experiment-specific parameters of safety index design for PPO-ISSA.
@@
8.3. Evaluating PPO-ISSA and Comparison Analysis
@@
Figure 11. Cost changes over 100 time steps of PPO and PPO-ISSA-CTrigger starting from the same unsafe state over 20 trails.
PPO-Lagrangian and CPO fail since both methods rely on trial-and-error to enforce constraints while ISSA is able to guarantee forward invariance by Theorem 1. We also observe that PPO-SL fails to lower the violation during training, due to the fact that the linear approximation of cost function $c(x_{t+1})\approx c(x_{t})+g(x_{t},w)^{T}u$ (dalal2018safe, 10) becomes inaccurate when the dynamics are highly nonlinear like the ones we used in MuJoCo (todorov2012mujoco, 37). More importantly, PPO-SL cannot guarantee that these always exist a feasible safe control to lower the cost, since they directly use the user defined cost function which cannot always ensure feasibility. More detailed metrics for comparison and experimental results on test suites with 1 constraint are summarized in Section 8.3.1.
8.3.1. Metrics Comparison
In this section, we report all the results of eight test suites by three metrics defined in Safety Gym (ray2019benchmarking, 32):
@@ 8.3.1. Metrics Comparison
The average episode return $J_{r}$ and the average episodic sum of costs $M_{c}$ were obtained by averaging over the last five epochs of training to reduce noise. Cost rate $\rho_{c}$ was just taken from the final epoch. We report the results of these three metrics in Table 4(h) normalized by PPO results. We calculate the converged reward $\bar{J}_{r}$ percentage of PPO-ISSA compared to other three safe RL baseline methods (PPO-Lagrangian, CPO and PPO-SL) over eight control suites. The computed mean reward percentage is $95\%$, and the standard deviation is $9\%$. Therefore we conclude that PPO-ISSA is able to gain $95\%\pm 9\%$ cumulative reward compared to state-of-the-art safe DRL methods.
@@
Table 4. Normalized metrics obtained from the policies at the end of the training process, which is averaged over eight test suits environments and five random seeds.
@@
Figure 13. Average episodic return, episodic cost and overall cost rate of constraints of PPO-ISSA and baseline methods on 4-constraint environments over five seeds.
@@
Figure 15. Average performance of PPO-ISSA and baseline methods on 1-constraint environments over five random seeds. The three rows represent average episodic return, average episodic cost and overall cost rate of constraints. The safety advantage of safe RL baseline methods over unconstrained RL method (PPO) becomes trivial as the constraint number and constraint size decrease, where the cost rate and average episode cost of PPO-Lagrangian, CPO and PPO are nearly the same when there is only one constraint with size 0.05.
Figure 16. Average return of PPO-ISSA with different safety index design on Goal-Hazard4-0.15.
To validate the finite time convergence of Theorem 1 that ISSA combined with CTrigger algorithm (PPO-ISSA-CTrigger) can ensure the finite time convergence to $\mathcal{X}_{S}$ for discrete-time system, we further compare the cost evolution of PPO and PPO-ISSA-CTrigger agents when starting from the same unsafe state (i.e., cost >0>0). The comparison results are shown in Figure 11, which shows that PPO-ISSA-CTrigger can converge to $\mathcal{X}_{S}$ within 100 time steps across all experiments while cost evolution of PPO agents fluctuates wildly without preference to converge to safe set. The cost changes of PPO-ISSA-CTrigger aligns with ours theory of finite time convergence.
8.4. Feasibility of Safety Index Synthesis
@@
Figure 17. Heat maps of the difference of safety index $\Delta\phi=\phi(f(x,u))-\phi(x)$. The x-axis $u_{1}$ represents the control space of moving actuator, and the y-axis $u_{2}$ represents the control space of turning actuator.
@@
8.5. Ablation Study
8.5.1. Sensitivity Analysis
@@
Table 5. Normalized computation time and return under different number of vectors in ISSA. These results are average on 100 ISSA runs over five random seeds on Goal-Hazard4-0.15.
@@
Figure 18. Average episodic return and episodic cost of PPO-ISSA and baseline methods on Goal-hazard1-0.15 environment of a doggo robot over five seeds.
@@ 8.5.2. Scalability Analysis
Figure 19. doggo robot: a quadrupedal robot with bilateral symmetry with 12-dimensional control space.
@@
Figure 20. Average performance of PPO-ISSA and baseline methods on Goal-hazard1-0.15 environment of doggo robot over five seeds. The three columns represent average episode return, average episode cost and overallcost rate of constraints.
8.5.3. ISSA in higher dimensional constrol systems
We test ISSA with a doggo robot as shown in Figure 19 with 12 dimensional control space and 80 dimensional state space. We evaluate ISSA with the doggo robot in the Goal-Hazard1-0.15 suite. As shown in Figure 20, ISSA is able to guarantee zero-violation in higher dimensional control systems. We notice that in system with higher control dimensions, safe RL methods like PPO-Lagrangian and CPO perform poorly compared to PPO, which demonstrates the constrained RL algorithms struggle to learn good reward performance for complex locomotion behavior. Similar comparison results are also reported in Safety Gym benchmarks (ray2019benchmarking, 32). In contrast, PPO-ISSA is able to achieve the best reward performance while guaranteeing zero safety violation, showing the scalability of ISSA to achieve satisfying reward performance in systems with higher control dimensions.
@@ 8.5.3. ISSA in higher dimensional constrol systems
Table 6. Average Computation time, succuss ratio of ISSA phase 1 and number of safe control candidates of 200 ISSA runs on Goal-Hazard1-0.15.
@@
8.5.4. AdamBA in higher dimensional space
@@
Figure 21. The average computation time of non-parallel AdamBA and parallel AdamBA on toy problem with different control dimensions and number of vectors.
9. Conclusion and Future Work
@@
There are three major directions for future work. Firstly, we will further generalize the safety index synthesis rule to cover a wider range of applications other than collision avoidance in 2D. Secondly, we will further speed up the implicit model evaluation step by replacing the physical engine based simulator with a learned surrogate model while taking the learned dynamics error into account. Third, obstacle avoidance tasks in densely cluttered environments — where multiple safety-critical constraints may be active simultaneously — still require further theoretical advancements and algorithmic exploration.
Acknowledgements.
This work is in part supported by Amazon Research Award and National Science Foundation through Grant #2144489. We also thank Tairan He for assistance with experiments.
Appendix A Reproducibility Checklist for JAIR
Select the answers that apply to your research – one per item.
All articles:
  1. (1)
    All claims investigated in this work are clearly stated. [yes]
  2. (2)
    Clear explanations are given how the work reported substantiates the claims. [yes]
  3. (3)
    Limitations or technical assumptions are stated clearly and explicitly. [yes]
  4. (4)
    Conceptual outlines and/or pseudo-code descriptions of the AI methods introduced in this work are provided, and important implementation details are discussed. [yes]
  5. (5)
    Motivation is provided for all design choices, including algorithms, implementation choices, parameters, data sets and experimental protocols beyond metrics. [yes]
Articles containing theoretical contributions:
Does this paper make theoretical contributions? [yes]
If yes, please complete the list below.
  1. (1)
    All assumptions and restrictions are stated clearly and formally. [yes]
  2. (2)
    All novel claims are stated formally (e.g., in theorem statements). [yes]
  3. (3)
    Proofs of all non-trivial claims are provided in sufficient detail to permit verification by readers with a reasonable degree of expertise (e.g., that expected from a PhD candidate in the same area of AI). [yes]
  4. (4)
    Complex formalism, such as definitions or proofs, is motivated and explained clearly. [yes]
  5. (5)
    The use of mathematical notation and formalism serves the purpose of enhancing clarity and precision; gratuitous use of mathematical formalism (i.e., use that does not enhance clarity or precision) is avoided. [yes]
  6. (6)
    Appropriate citations are given for all non-trivial theoretical tools and techniques. [yes]
Articles reporting on computational experiments:
Does this paper include computational experiments? [yes]
If yes, please complete the list below.
  1. (1)
    All source code required for conducting experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for source code readability and documentation as well as for long-term accessibility. [yes]
  2. (2)
    The source code comes with a license that allows free usage for reproducibility purposes. [yes]
  3. (3)
    The source code comes with a license that allows free usage for research purposes in general. [yes]
  4. (4)
    Raw, unaggregated data from all experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility. [yes]
  5. (5)
    The unaggregated data comes with a license that allows free usage for reproducibility purposes. [yes]
  6. (6)
    The unaggregated data comes with a license that allows free usage for research purposes in general. [yes]
  7. (7)
    If an algorithm depends on randomness, then the method used for generating random numbers and for setting seeds is described in a way sufficient to allow replication of results. [yes]
  8. (8)
    The execution environment for experiments, the computing infrastructure (hardware and software) used for running them, is described, including GPU/CPU makes and models; amount of memory (cache and RAM); make and version of operating system; names and versions of relevant software libraries and frameworks. [yes]
  9. (9)
    The evaluation metrics used in experiments are clearly explained and their choice is explicitly motivated. [yes]
  10. (10)
    The number of algorithm runs used to compute each result is reported. [yes]
  11. (11)
    Reported results have not been “cherry-picked” by silently ignoring unsuccessful or unsatisfactory experiments. [yes]
  12. (12)
    Analysis of results goes beyond single-dimensional summaries of performance (e.g., average, median) to include measures of variation, confidence, or other distributional information. [yes]
  13. (13)
    All (hyper-) parameter settings for the algorithms/methods used in experiments have been reported, along with the rationale or method for determining them. [yes]
  14. (14)
    The number and range of (hyper-) parameter settings explored prior to conducting final experiments have been indicated, along with the effort spent on (hyper-) parameter optimisation. [yes]
  15. (15)
    Appropriately chosen statistical hypothesis tests are used to establish statistical significance in the presence of noise effects. [yes]
Articles using data sets:
Does this work rely on one or more data sets (possibly obtained from a benchmark generator or similar software artifact)? [no]
If yes, please complete the list below.
  1. (1)
    All newly introduced data sets are included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility with a license that allows free usage for research purposes. [yes/partially/no/NA]
  2. (2)
    The newly introduced data set comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
  3. (3)
    The newly introduced data set comes with a license that allows free usage for research purposes in general. [yes/partially/no]
  4. (4)
    All data sets drawn from the literature or other public sources (potentially including authors’ own previously published work) are accompanied by appropriate citations. [yes/no/NA]
  5. (5)
    All data sets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. [yes/partially/no/NA]
  6. (6)
    All new data sets and data sets that are not publicly available are described in detail, including relevant statistics, the data collection process and annotation process if relevant. [yes/partially/no/NA]
  7. (7)
    All methods used for preprocessing, augmenting, batching or splitting data sets (e.g., in the context of hold-out or cross-validation) are described in detail. [yes/partially/no/NA]